A maximizing worker with u(z, f) = z1/2f1/2 earns 60 dollars per day in non-labour income and is paid a wage of 5 dollars per hour. How many hours will she work. If she joins a union she will be paid a wage of 20 dollars per hour but she will have to pay a membership fee. What is the maximum daily fee she would pay to join this union?



1
Expert's answer
2020-10-29T07:05:47-0400

The reservation wage is: rw = 60/24 = 2.5. So, she will decide to work and will work at least 8 hours per day. If she joins a union, then the maximum daily fee she would pay to join this union is (20 - 5)×8 = $120.